# Discovery and Analysis of Afterglows from Poorly Localised GRBs with the Gravitational-wave Optical Transient Observer (GOTO) All-sky Survey

### v1

@@ 5.3 $T_{90}$ comparison
Table 5: Summary of Swift/XRT afterglow spectral and temporal properties for our GRB sample. Exp. denotes the spectral extraction time interval. $N_{\rm H,Gal}$ is fixed to the Galactic line-of-sight value, while $N_{\rm H,int}$ is the intrinsic absorption component derived from spectral fitting. Fluxes are unabsorbed values in the 0.3–10.0 keV band. Negative $\alpha$ values indicate apparent rising trends likely caused by low-count statistics.
@@ 6.1 X-ray
Spectra for both source and background regions, along with corresponding ancillary and response files extracted in PC mode from the Swift/XRT repository, were grouped to a minimum of one count per bin using grppha task. Spectral fitting was performed in XSPEC using Cash statistics (Cash, 1979), appropriate for low-count Poisson-distributed data. Each spectrum was modelled with a simple absorbed powerlaw, tbabs*ztbabs(zpowerlw) (Wilms et al., 2000), accounting for both Galactic foreground and intrinsic absorption in the host galaxy. Galactic column density ($N_{\rm H,g}$) for each GRB was fixed at a value obtained from Swift Galactic $N_{\rm H}$ tool [^17] https://www.swift.ac.uk/analysis/nhtot/ (Willingale et al., 2013), while the intrinsic host galactic absorption ($N_{\rm H,intr}$) and photon spectral index ($\Gamma$) were left free to vary. All fits were performed in the 0.3–10.0 keV energy range, and the errors reported here correspond to 90% confidence intervals. Figure A12 presents the 0.3–10.0 keV XRT count rate spectra for all seven GRBs in our sample, overlaid (red solid line) with their respective best-fit absorbed powerlaw models.
Table 5 summarises the temporal and spectral properties derived from Swift/XRT observations of our GRB sample. Redshifts used for spectral fitting are reported in later sections. Photon indices lie in the range $\Gamma\simeq 1.4-2.9$, consistent with typical afterglow spectra. The intrinsic absorption shows substantial variation: GRBs 240122A and 240910A are consistent with negligible additional absorption beyond the Galactic foreground, whereas others, most notably GRB 240916A, require higher column densities of the order of $10^{22}$ cm-2. The inferred unabsorbed 0.3–10 keV fluxes span nearly two orders of magnitude, from $\sim 7\times 10^{-13}$ to $\sim 1.3\times 10^{-11}$ erg cm-2 s-1, with the highest values observed for GRBs 240225B and 240122A, despite their relatively short spectral extraction intervals. Temporal decay slopes cluster around $\alpha\approx 1.0-1.3$ for most afterglows, though three cases (GRBs 240619A, 240910A, and 241228B) show formally flat or rising indices; in each case, the uncertainties are large and the behaviour is consistent with constant flux within errors. At such late phases, exposure times were modest (typically 1–3 ks), naturally limiting the statistical precision of the spectral fits.
@@ 6.3.3 Redshift Comparison
Figure 14: Multi-wavelength afterglow light curves for six GRBs modelled in this work: GRBs 240122A, 240225B, 240619A, 240910A, 240916A, and 241228B. Each panel shows the observed data points (UV/optical/NIR, X-ray, and radio, where available), overlaid with the best-fit afterglow model using a TopHat (uniform) jet scenario (see text for details). Photometric data are compiled from our observations and published GCN Circulars. Fluxes are rescaled for clarity as indicated in the legends.
@@
Figure 15: Posterior distributions for the TopHat-jet afterglow model parameters inferred for six GRBs analysed in this work: GRBs 240122A, 240225B, 240619A, 240910A, 240916A, and 241228B. Each panel shows the marginalised 1D and 2D posterior distributions from the dynesty nested-sampling run (via Bilby), with contours representing $68\%$ and $95\%$ credible regions. Inferred parameters include the observer angle, core angle, isotropic-equivalent energy, and microphysical quantities (see text for details).
@@ 7 Afterglow modelling
Overall, the parameters inferred from our TopHat (uniform) jet fits — specifically the observer angle ($\theta_{v}$), the jet half-opening angle ($\theta_{j}\equiv\theta_{c}$ in this model), and the isotropic-equivalent kinetic energy ($E_{0}$) — are consistent with the ranges reported in previous studies of long-duration GRB afterglows. In all cases, we recover $\theta_{v}$ and $\theta_{j}$ of order a few degrees, in line with the narrow jet geometries commonly found in broadband afterglow fits (e.g. Frail et al. (2001); Bloom et al. (2003); Friedman & Bloom (2005); Racusin et al. (2009)). We use $E_{0}$ to estimate the beaming-corrected jet kinetic energy, yielding values that cluster around $10^{51}-10^{52}$ erg, consistent with the canonical long-GRB energy scale (e.g. Friedman & Bloom (2005)). This agreement in both angular geometry and energetics supports the robustness of our TopHat-jet modelling and places these events firmly within the established population of classical long-duration GRBs.
@@ Acknowledgements
It is also partly based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. The NOT data were obtained under program IDs 68-020, 69-023, and 70-507 (PIs Malesani, Fynbo, Xu). The National Radio Astronomy Observatory and Green Bank Observatory are facilities of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for World-Class University" and “Yunnan University Development Plan for World-Class Astronomy Discipline". Based on observations made at the Observatorio de Sierra Nevada (OSN), operated by the Instituto de Astrofísica de Andalucía (IAA-CSIC).
@@
AK and JRM are supported by the UK Science and Technology Facilities Council (STFC) Consolidated grant ST/V000853/1. BPG acknowledges support from STFC grant No. ST/Y002253/1. BPG and DO acknowledge support from The Leverhulme Trust grant No. RPG-2024-117. BS acknowledges the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE31-0011 (project PEGaSUS). DS is supported by the UK Science and Technology Facilities Council (STFC, grant numbers ST/T007184/1, ST/T003103/1, and ST/T000406/1). SG acknowledges support from the Istituto Nazionale di Astrofisica (INAF), project number: 1.05.24.07.04. BK acknowledges support from the “Special Project for High-End Foreign Experts," Xingdian Funding from Yunnan Province; the Yunnan Key Laboratory of Survey Science (grant 202449CE340002), and National Key Research and Development Program of China (grant 2024YFA1611603). BW is supported by the UKRI’s STFC studentship grant funding, project reference ST/X508871/1 RLCS and SM acknowledge support from The Leverhulme Trust grant RPG-2023-240. NH, POB and NRT acknowledge support from UKRI/STFC grant ST/W000857/1. JDL acknowledges support from a UK Research and Innovation Future Leaders Fellowship (MR/T020784/1). TLK acknowledges a Warwick Astrophysics prize post-doctoral fellowship made possible thanks to a generous philanthropic donation. BS and SDV acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE31-0011 (project PEGaSUS). MAA acknowledges support from grants PID2021-127495NB-I00 (funded by MCIN/AEI/10.13039/501100011033, EU), ASFAE/2022/026 (funded by MCIN, EU NextGenerationEU PRTR-C17.I1), and CIPROM/2022/13 (Generalitat Valenciana). DLC acknowledges support from the UK Science and Technology Facilities Council (STFC) grant number ST/X001121/1. AS acknowledges support by a postdoctoral fellowship from the CNES. AR acknowledges support from the INAF project Supporto Arizona & Italia. AMC and LC acknowledge support from the Irish Research Council Postgraduate Scholarship No. GOIPG/2022/1008. MEW and IW are supported by the UKRI Science and Technology Facilities Council (STFC) DBM is funded by the European Union (ERC, HEAVYMETAL, 101071865). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. NSP and ASP are grateful to the Russian Science Foundation (project no. 23-12-00220) for their partial support of the data reduction, analysis of data. EVK is grateful to the Ministry of Science and Higher Education of the Russian Federation for financial support of this work; the AZT-33IK telescope is part of the Center for Common Use "Angara".
@@
Data can be shared upon request to the corresponding author.
@@
Figure A9: MAXI fitted spectra for GRBs 240122A and 240225B are shown. Models shown are the best-fit models as described in Table 4. GRB 241228B is poorly constrained in high energies, so the spectrum shown is zoomed in to show the behaviour at lower energies.
@@
Figure A11: Fermi/GBM fitted spectra for Fermi GRBs in our sample are shown. Models shown are the best-fit models as described in Table 4. GRB 241228B is poorly constrained in high energies, so the spectrum shown is zoomed in to show the behaviour at lower energies.
Figure A12: Swift/XRT count-rate spectra and best-fitting X-ray continuum models for the GRBs in our sample (GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B). In each panel, the top plot displays observed spectral data with errors (black points) in the 0.3–10 keV range fitted with an absorbed power-law model (red line), using Cash statistics. The bottom plot shows the ratio of observed data to folded model predictions, used to assess the goodness of fit. These spectra were used to derive photon indices and absorption column densities ($N_{\mathrm{H}}$), contributing to the multi-wavelength characterisation of the afterglows.
@@
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
@@
Table A5: Absorption and emission features identified in the afterglow spectrum of GRB 240619A ($z=0.3960\pm 0.0001$), obtained with VLT/X-shooter.
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type
@@
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ [Å] $\lambda_{\textrm{rest}}$ [Å] Feature Type Arm
@@
1Department of Physics, Royal Holloway - University of London, Egham, TW20 0EX, U.K.
@@
2Department of Physics, University of Warwick, Gibbet Hill Road, Coventry CV4 7AL, UK.
3School of Physics and Astronomy, University of Birmingham, Edgbaston, Birmingham, B15 2TT, UK.
4Institute for Gravitational Wave Astronomy, University of Birmingham, Birmingham, B15 2TT, UK.
5Aix Marseille Univ., CNRS, CNES, LAM, Marseille, France.
6School of Physics & Astronomy, Monash University, Clayton VIC 3800, Australia.
7Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, 91191, Gif-sur-Yvette, France.
8School of Physics and Astronomy, University of Leicester, University Road, Leicester, LE1 7RH, United Kingdom.
9Sydney Institute for Astronomy, School of Physics, The University of Sydney, NSW 2006, Australia.
@@
10Centre of Excellence for Gravitational Wave Discovery (OzGrav), Hawthorn, VIC 3122, Australia.
11CSIRO Space and Astronomy, PO Box 76, Epping, NSW 1710, Australia.
12Department of Astrophysics/IMAPP, Radboud University, 6525 AJ Nijmegen, The Netherlands.
13Niels Bohr Institute, University of Copenhagen, Jagtvej 155, 2200, Copenhagen N, Denmark
14The Cosmic Dawn Centre (DAWN), Denmark
@@
15Astrophysics Research Cluster, School of Mathematical and Physical Sciences, University of Sheffield, Sheffield, S3 7RH, UK.
16Research Software Engineering, University of Sheffield, Sheffield, S1 4DP, UK.
17INAF Osservatorio Astronomico di Brera, Via E. Bianchi 46, I-23807 Merate, Italy.
@@
18Department of Physical Science, Aoyama Gakuin University, 5-10-1 Fuchinobe, Chuo-ku, Sagamihara, Kanagawa 252-5258, Japan.
@@
19South-Western Institute for Astronomy Research (SWIFAR), Yunnan University, Kunming, Yunnan 650500, People’s Republic of China.
@@
20Yunnan Key Laboratory of Survey Science, Yunnan University, Kunming, Yunnan 650500, People’s Republic of China.
@@
21National Astronomical Observatories, Chinese Academy of Sciences, Beijing 100012, China.
@@
22Guangxi Key Laboratory for Relativistic Astrophysics, School of Physical Science and Technology, Guangxi University, Nanning 530004, China.
23Armagh Observatory and Planetarium, College Hill, Armagh, BT61 9DG, N Ireland, UK.
24Institute for Globally Distributed Open Research and Education (IGDORE).
25Instituto de Astrofísica de Canarias, E-38205 La Laguna, Tenerife, Spain.
@@
26National Astronomical Research Institute of Thailand, 260 Moo 4, T. Donkaew, A. Maerim, Chiangmai, 50180 Thailand.
27Department of Physics & Astronomy, University of Turku, Vesilinnantie 5, Turku, FI-20014, Finland.
@@
28Jodrell Bank Centre for Astrophysics, Department of Physics and Astronomy, The University of Manchester, Manchester, M13 9PL, UK.
29Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK.
30School of Physics, University College Cork, Cork, Ireland.
@@
31Centre for Astrophysics Research, University of Hertfordshire, College Lane, Hatfield, AL10 9AB, UK.
@@
32School of Physics and Centre for Space Research, University College Dublin, Belfield, Dublin 4, Ireland.
@@
33National Research University ‘Higher School of Economics’, Faculty of Physics, Myasnitskaya ul. 20, Moscow 101000, Russia.
@@
34Space Research Institute of the Russian Academy of Sciences, Profsoyuznaya ul. 84/32, Moscow 117997, Russia.
@@
35Centro Astronómico Hispano en Andalucía, Observatorio de Calar Alto, Sierra de los Filabres, Gérgal, Almería, 04550, Spain
@@
36Departament d’Astronomía i Astrofísica, Universitat de València, 46100 Burjassot, Spain
37Observatori Astronòmic, Universitat de València, 46980 Paterna, Spain
@@
38International Centre for Radio Astronomy Research, Curtin University, GPO Box U1987, Perth, WA 6845, Australia.
@@
39 Astrophysics Research Institute, Liverpool John Moores University, 146 Brownlow Hill, Liverpool L3 5RF, UK
@@
40Instituto de Astrofísica de Andalucía (IAA-CSIC), Glorieta de la Astronomía s/n, E-18008, Granada, Spain.
@@
41Ingeniería de Sistemas y Automática, Universidad de Málaga, Unidad Asociada al CSIC porel IAA, Escuela de Ingenierías Industriales, Arquitecto Francisco Peñalosa, 6, Campanillas, 29071 Málaga, Spain.
@@
42Millennium Institute of Astrophysics (MAS), Nuncio Monsenor Sòtero Sanz 100, Providencia, Santiago RM, Chile.
43University of Messina, MIFT Department, Via F.S. D’Alcontres 31, 98166, Messina, Italy
@@
44Space Science Data Center (SSDC) - Agenzia Spaziale Italiana (ASI), Via del Politecnico snc, I-00133 Roma, Italy
@@
45Department of Astronomy, School of Physics, Huazhong University of Science and Technology, Wuhan, 430074, China
46Clemson University, Department of Physics and Astronomy, Clemson, SC 29634-0978, USA
@@
47Osservatorio Astronomico di Capodimonte, INAF, Salita Moiariello 16, Napoli, 80131, Italy
@@
48Institute of Solar-Terrestrial Physics, Russian Academy of Sciences, Siberian Branch, Irkutsk 664033, Russia.
49GRANTECAN S.A., Cuesta de San José s/n, E-38712 Breña Baja, La Palma, Spain.
@@
50INAF - Osservatorio di Astrofisica e Scienza dello Spazio, via Piero Gobetti 93/3, 40129 Bologna, Italy

### v2

@@ 5.3 $T_{90}$ comparison
Table 5: Summary of Swift/XRT afterglow spectral and temporal properties for our GRB sample. Exp. denotes the spectral extraction time interval. $N_{\rm H,Gal}$ is fixed to the Galactic line-of-sight value, while $N_{\rm H,int}$ is the intrinsic absorption component derived from spectral fitting. Fluxes are unabsorbed values in the 0.3–10.0 keV band. $\alpha_{X}$ is the X-ray temporal index, where negative values indicate apparent rising trends likely caused by low-count statistics.
@@ 6.1 X-ray
Spectra for both source and background regions, along with corresponding ancillary and response files extracted in PC mode from the Swift/XRT repository, were grouped to a minimum of one count per bin using grppha task. Spectral fitting was performed in XSPEC using Cash statistics (Cash, 1979), appropriate for low-count Poisson-distributed data. Each spectrum was modelled with a simple absorbed powerlaw, tbabs*ztbabs(zpowerlw) (Wilms et al., 2000), accounting for both Galactic foreground and intrinsic absorption in the host galaxy. Galactic column density ($N_{\rm H,g}$) for each GRB was fixed at a value obtained from Swift Galactic $N_{\rm H}$ tool [^17] https://www.swift.ac.uk/analysis/nhtot/ (Willingale et al., 2013), while the intrinsic host galactic absorption ($N_{\rm H,intr}$) and photon index ($\Gamma$) were left free to vary. All fits were performed in the 0.3–10.0 keV energy range, and the errors reported here correspond to 90% confidence intervals. Figure A12 presents the 0.3–10.0 keV XRT count rate spectra for all seven GRBs in our sample, overlaid (red solid line) with their respective best-fitting absorbed powerlaw models.
Table 5 summarises the temporal and spectral properties derived from Swift/XRT observations of our GRB sample. Redshifts used for spectral fitting are reported in later sections. Photon indices lie in the range $\Gamma\simeq 1.4-2.9$, consistent with typical afterglow spectra. The intrinsic absorption shows substantial variation: GRBs 240122A and 240910A are consistent with negligible additional absorption beyond the Galactic foreground, whereas others, most notably GRB 240916A, require higher column densities of the order of $10^{22}$ cm-2. The inferred unabsorbed 0.3–10 keV fluxes span nearly two orders of magnitude, from $\sim 7\times 10^{-13}$ to $\sim 1.3\times 10^{-11}$ erg cm-2 s-1, with the highest values observed for GRBs 240225B and 240122A, despite their relatively short spectral extraction intervals. Temporal decay slopes cluster around $\alpha_{X}\approx 1.0-1.3$ for most afterglows, though three cases (GRBs 240619A, 240910A, and 241228B) show formally flat or rising indices; in each case, the uncertainties are large and the behaviour is consistent with constant flux within errors. At such late phases, exposure times were modest (typically 1–3 ks), naturally limiting the statistical precision of the spectral fits.
@@ 6.3.3 Redshift Comparison
Figure 14: Multi-wavelength afterglow light curves for six GRBs modelled in this work: GRBs 240122A, 240225B, 240619A, 240910A, 240916A, and 241228B. Each panel shows the observed data points (UV/optical/NIR, X-ray, and radio, where available), overlaid with the best-fitting afterglow model using a TopHat (uniform) jet scenario (see the text for details). Photometric data are compiled from our observations and published GCN Circulars. Fluxes are rescaled for clarity as indicated in the legends.
@@
Figure 15: Posterior distributions for the TopHat-jet afterglow model parameters inferred for six GRBs analysed in this work: GRBs 240122A, 240225B, 240619A, 240910A, 240916A, and 241228B. Each panel shows the marginalised 1D and 2D posterior distributions from the dynesty nested-sampling run (via Bilby), with contours representing $68\%$ and $95\%$ credible regions. Inferred parameters include the observer angle, core angle, isotropic-equivalent energy, and microphysical quantities (see the text for details).
@@ 7 Afterglow modelling
Overall, the parameters inferred from our TopHat (uniform) jet fits — specifically the observer angle ($\theta_{v}$), the jet half-opening angle ($\theta_{j}\equiv\theta_{c}$ in this model), and the isotropic-equivalent kinetic energy ($E_{0}$) — are consistent with the ranges reported in previous studies of long-duration GRB afterglows. In all cases, we recover $\theta_{v}$ and $\theta_{j}$ of order a few degrees, in line with the narrow jet geometries commonly found in broadband afterglow fits (e.g. Frail et al. 2001; Bloom et al. 2003; Friedman & Bloom 2005; Racusin et al. 2009). We use $E_{0}$ to estimate the beaming-corrected jet kinetic energy, yielding values that cluster around $10^{51}-10^{52}$ erg, consistent with the canonical long-GRB energy scale (e.g. Friedman & Bloom (2005)). This agreement in both angular geometry and energetics supports the robustness of our TopHat-jet modelling and places these events firmly within the established population of classical long-duration GRBs.
@@ Acknowledgements
It is also partly based on data obtained with the ATCA and we acknowledge the Gomeroi people as the traditional owners of the ATCA observatory site. The ATCA is part of the Australia Telescope National Facility, which is funded by the Australian Government for operation as a National Facility managed by CSIRO.
@@
It is also partly based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. The NOT data were obtained under program IDs 68-020, 69-023, and 70-507 (PIs Malesani, Fynbo, Xu). The National Radio Astronomy Observatory and Green Bank Observatory are facilities of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. Mephisto is developed at and operated by the South-Western Institute for Astronomy Research of Yunnan University (SWIFAR-YNU), funded by the “Yunnan University Development Plan for World-Class University" and “Yunnan University Development Plan for World-Class Astronomy Discipline". Based on observations made at the Observatorio de Sierra Nevada (OSN), operated by the Instituto de Astrofísica de Andalucía (IAA-CSIC). Based on observations made with the Thai Robotic Telescopes under program ID TRTToO_2024005, which is operated by the National Astronomical Research Institute of Thailand (Public Organisation).
@@
AK and JRM are supported by the UK Science and Technology Facilities Council (STFC) Consolidated grant ST/V000853/1. BPG acknowledges support from STFC grant No. ST/Y002253/1. BPG and DO acknowledge support from The Leverhulme Trust grant No. RPG-2024-117. BS and SDV acknowledge the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-23-CE31-0011 (project PEGaSUS). DS is supported by the UK Science and Technology Facilities Council (STFC, grant numbers ST/T007184/1, ST/T003103/1, and ST/T000406/1). SG acknowledges support from the Istituto Nazionale di Astrofisica (INAF), project number: 1.05.24.07.04. BK acknowledges support from the “Special Project for High-End Foreign Experts," Xingdian Funding from Yunnan Province; the Yunnan Key Laboratory of Survey Science (grant 202449CE340002), and National Key Research and Development Program of China (grant 2024YFA1611603). BW is supported by the UKRI’s STFC studentship grant funding, project reference ST/X508871/1. RLCS and SM acknowledge support from The Leverhulme Trust grant RPG-2023-240. NH, POB and NRT acknowledge support from UKRI/STFC grant ST/W000857/1. JDL acknowledges support from a UK Research and Innovation Future Leaders Fellowship (MR/T020784/1). TLK acknowledges a Warwick Astrophysics prize post-doctoral fellowship made possible thanks to a generous philanthropic donation. MAA acknowledges support from grants PID2021-127495NB-I00 (funded by MCIN/AEI/10.13039/501100011033, EU), ASFAE/2022/026 (funded by MCIN, EU NextGenerationEU PRTR-C17.I1), and CIPROM/2022/13 (Generalitat Valenciana). DLC acknowledges support from the UK Science and Technology Facilities Council (STFC) grant number ST/X001121/1. AS acknowledges support by a postdoctoral fellowship from the CNES. AR acknowledges support from the INAF project Supporto Arizona & Italia. AMC and LC acknowledge support from the Irish Research Council Postgraduate Scholarship No. GOIPG/2022/1008. MEW and IW are supported by the UKRI Science and Technology Facilities Council (STFC). DBM is funded by the European Union (ERC, HEAVYMETAL, 101071865). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. NSP and ASP are grateful to the Russian Science Foundation (project no. 23-12-00220) for their partial support of the data reduction, analysis of data. EVK is grateful to the Ministry of Science and Higher Education of the Russian Federation for financial support of this work; the AZT-33IK telescope is part of the Center for Common Use "Angara".
@@
All datasets supporting this study are provided in the paper. Additional data are available from the corresponding author upon reasonable request.
@@
Figure A9: MAXI fitted spectra for GRBs 240122A and 240225B are shown. Models shown are the best-fitting models as described in Table 4.
@@
Figure A11: Fermi/GBM fitted spectra for Fermi GRBs in our sample are shown. Models shown are the best-fitting models as described in Table 4. GRB 241228B is poorly constrained in high energies, so the spectrum shown is zoomed in to show the behaviour at lower energies.
Figure A12: Swift/XRT count-rate spectra and best-fitting X-ray continuum models for the GRBs in our sample. In each panel, the top plot displays observed spectral data with errors (black points) in the 0.3–10 keV range fitted with an absorbed power-law model (red line), using Cash statistics. The bottom plot shows the ratio of observed data to folded model predictions, used to assess the goodness of fit. These spectra were used to derive photon indices and absorption column densities ($N_{\mathrm{H}}$), contributing to the multi-wavelength characterisation of the afterglows.
@@
$\lambda_{\textrm{obs}}$ (Å) $\lambda_{\textrm{rest}}$ (Å) Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ (Å) $\lambda_{\textrm{rest}}$ (Å) Feature Type Arm
@@
Table A5: Absorption and emission features identified in the afterglow spectrum of GRB 240619A ($z=0.3960\pm 0.0001$), obtained with VLT/X-shooter. Here, a marks lines affected by telluric absorption.
$\lambda_{\textrm{obs}}$ (Å) $\lambda_{\textrm{rest}}$ (Å) Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ (Å) $\lambda_{\textrm{rest}}$ (Å) Feature Type
@@
$\lambda_{\textrm{obs}}$ (Å) $\lambda_{\textrm{rest}}$ (Å) Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ (Å) $\lambda_{\textrm{rest}}$ (Å) Feature Type Arm
@@
$\lambda_{\textrm{obs}}$ (Å) $\lambda_{\textrm{rest}}$ (Å) Feature Type Arm
@@
1Department of Physics, Royal Holloway - University of London, Egham, TW20 0EX, UK
@@
2Department of Physics, University of Warwick, Gibbet Hill Road, Coventry CV4 7AL, UK
3School of Physics and Astronomy, University of Birmingham, Edgbaston, Birmingham, B15 2TT, UK
4Institute for Gravitational Wave Astronomy, University of Birmingham, Birmingham, B15 2TT, UK
5Aix Marseille Univ., CNRS, CNES, LAM, Marseille, F-13388, France
6School of Physics & Astronomy, Monash University, Clayton VIC 3800, Australia
7Université Paris-Saclay, Université Paris Cité, CEA, CNRS, AIM, 91191, Gif-sur-Yvette, France
8School of Physics and Astronomy, University of Leicester, University Road, Leicester, LE1 7RH, UK
9Sydney Institute for Astronomy, School of Physics, The University of Sydney, NSW 2006, Australia
@@
10Centre of Excellence for Gravitational Wave Discovery (OzGrav), Hawthorn, VIC 3122, Australia
11CSIRO Space and Astronomy, PO Box 76, Epping, NSW 1710, Australia
12The Cosmic Dawn Centre (DAWN), Rådmandsgade 64, DK-2200, København N., Denmark
13Niels Bohr Institute, University of Copenhagen, Jagtvej 155, DK-2200 Copenhagen N, Denmark
14Department of Astrophysics/IMAPP, Radboud University, NL-6525 AJ Nijmegen, the Netherlands
@@
15Astrophysics Research Cluster, School of Mathematical and Physical Sciences, University of Sheffield, Sheffield, S3 7RH, UK
16Research Software Engineering, University of Sheffield, Sheffield, S1 4DP, UK
17INAF Osservatorio Astronomico di Brera, Via E. Bianchi 46, I-23807 Merate, Italy
@@
18Department of Physical Science, Aoyama Gakuin University, 5-10-1 Fuchinobe, Chuo-ku, Sagamihara, Kanagawa 252-5258, Japan
@@
19South-Western Institute for Astronomy Research (SWIFAR), Yunnan University, Kunming, Yunnan 650500, People’s Republic of China
@@
20Yunnan Key Laboratory of Survey Science, Yunnan University, Kunming, Yunnan 650500, People’s Republic of China
@@
21National Astronomical Observatories, Chinese Academy of Sciences, Beijing 100101, China
@@
22Guangxi Key Laboratory for Relativistic Astrophysics, School of Physical Science and Technology, Guangxi University, Nanning 530004, China
23Armagh Observatory and Planetarium, College Hill, Armagh, BT61 9DG, Northern Ireland, UK
24Institute for Globally Distributed Open Research and Education (IGDORE)
25Instituto de Astrofísica de Canarias, E-38205 La Laguna, Tenerife, Spain
@@
26National Astronomical Research Institute of Thailand, 260 Moo 4, T. Donkaew, A. Maerim, Chiangmai, 50180 Thailand
27Department of Physics & Astronomy, University of Turku, Vesilinnantie 5, Turku, FI-20014, Finland
@@
28Jodrell Bank Centre for Astrophysics, Department of Physics and Astronomy, The University of Manchester, Manchester, M13 9PL, UK
29Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK
30Departamento de Astrofísica, Universidad de La Laguna, E-38206 La Laguna, Tenerife, Spain
31School of Physics, University College Cork, Cork, T12 K8AF, Ireland
@@
32Centre for Astrophysics Research, University of Hertfordshire, College Lane, Hatfield AL10 9AB, UK
@@
33School of Physics and Centre for Space Research, University College Dublin, Belfield, Dublin 4, Ireland
@@
34National Research University ‘Higher School of Economics’, Faculty of Physics, Myasnitskaya ul. 20, Moscow 101000, Russia
@@
35Space Research Institute of the Russian Academy of Sciences, Profsoyuznaya ul. 84/32, Moscow 117997, Russia
@@
36Centro Astronómico Hispano en Andalucía, Observatorio de Calar Alto, Sierra de los Filabres, Gérgal, Almería, E-04550, Spain
@@
37Departament d’Astronomía i Astrofísica, Universitat de València, 46100 Burjassot, Spain
38Observatori Astronòmic, Universitat de València, 46980 Paterna, Spain
@@
39International Centre for Radio Astronomy Research, Curtin University, GPO Box U1987, Perth, WA 6845, Australia
@@
40Astrophysics Research Institute, Liverpool John Moores University, 146 Brownlow Hill, Liverpool L3 5RF, UK
@@
41Instituto de Astrofísica de Andalucía (IAA-CSIC), Glorieta de la Astronomía s/n, E-18008, Granada, Spain
@@
42Ingeniería de Sistemas y Automática, Universidad de Málaga, Unidad Asociada al CSIC por el IAA, Escuela de Ingenierías Industriales, Arquitecto Francisco Peñalosa, 6, Campanillas, 29071 Málaga, Spain
@@
43Millennium Institute of Astrophysics (MAS), Nuncio Monsenor Sòtero Sanz 100, Providencia, Santiago RM, 8320000, Chile
44MIFT Department, University of Messina, Via F.S. D’Alcontres 31, I-98166 Messina, Italy
@@
45Space Science Data Center (SSDC) – Agenzia Spaziale Italiana (ASI), Via del Politecnico snc, I-00133 Roma, Italy
@@
46Department of Astronomy, School of Physics, Huazhong University of Science and Technology, Wuhan 430074, China
47Clemson University, Department of Physics and Astronomy, Clemson, SC 29634-0978, USA
@@
48Osservatorio Astronomico di Capodimonte, INAF, Salita Moiariello 16, I-80131 Napoli, Italy
@@
49Institute of Solar-Terrestrial Physics, Russian Academy of Sciences, Siberian Branch, Irkutsk 664033, Russia
50GRANTECAN S.A., Cuesta de San José s/n, E-38712 Breña Baja, La Palma, Spain
@@
51INAF – Osservatorio di Astrofisica e Scienza dello Spazio, via Piero Gobetti 93/3, I-40129 Bologna, Italy